Of a plane takes off from an airport that sit at an elevation of 510 ft and ofter traveling 105 Miles has a flying altitude of 35,000 ft, what is the average slope of the plane ascent
Answer by Alan3354(69443)

Of a plane takes off from an airport that sit at an elevation of 510 ft and ofter traveling 105 Miles has a flying altitude of 35,000 ft, what is the average slope of the plane ascent
---------------------
Change in altitude (y) = 34490 feet
Change in distance (x) = 105*5280 = 554400 feet
---
tan = y/x = slope